OpenStudy (anonymous):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression.

OpenStudy (anonymous):

\[3\sqrt{2x} + x \sqrt{8x} - 5 \sqrt{18x} and got -10x \sqrt{2x}\] what did he do wrong?

OpenStudy (nory):

First: in all the terms, we need the number under the radical to be the same. √2x is the radical in the first term. Nothing to do there. √8x is the second radical. But 8 = 4*2, so we can write it as √8x = √4*2x = √4 * √2x = 2√2x. Now, what do you get for the third one?

OpenStudy (anonymous):

To be honest, I'm not good at these at all.

OpenStudy (anonymous):

I'm trying to understand what you just said. @Nory

OpenStudy (nory):

It's okay. I didn't explain it very well.

OpenStudy (nory):

So to add radicals, you have to get the same number under the radical sign in all the terms. Does that make any sense?

OpenStudy (anonymous):

Yes.

OpenStudy (nory):

But that is not the case in your problem...yet. We have to get the same number under the radical in all the terms. We can do this by finding a square number to factor out of the stuff under the radical. Ex. √4x well, we can just factor out the four... √4x = √4 * √x = 2√x Does that make sense? Maybe that's where you're getting stuck.

OpenStudy (anonymous):

Yes, that does make sense.

OpenStudy (nory):

Okay. Now we can do the same thing to our radicals. Let's work on x√8x. We can basically use the same technique as above. x√8x = x√4*2x = x√4 * √2x = 2x√2x Does that make sense?

OpenStudy (anonymous):

I see what you did there. I understand that part.

OpenStudy (nory):

Okay, good. Now 5√18x = 5√9 * √2x = 15√2x. That makes sense, right? Now write all our terms out in a line. 3√2x + 2x√2x - 15√2x. Does this look familiar?

OpenStudy (anonymous):

do you mean, \[10 \sqrt{18x}\] ?

OpenStudy (anonymous):

\[5\sqrt{9} * \sqrt{2}\]

OpenStudy (nory):

Go on...that √9 looks like it could be simplified.

OpenStudy (anonymous):

\[10\sqrt{3}\] ?

OpenStudy (anonymous):

@Nory

OpenStudy (nory):

I think I see where you got mixed up... 5√9 * √2 = 5*3 *√2 = 15√2

OpenStudy (anonymous):

Oh okay.

OpenStudy (nory):

Now you can add/subtract them as you would regular numbers.

OpenStudy (anonymous):

What about the 3 and the 2x?

OpenStudy (nory):

What 3? The 2x under the radical, you just tack on at the end. For example, to do 2√x + 5√x, you first do 2+5=7, then add the x. √7x.

OpenStudy (anonymous):

\[3\sqrt{2x} + 2x \sqrt{2x} - 15\sqrt{2x}\] You said it was this?

OpenStudy (nory):

Yeah, that looks like what we got. Go on...?

OpenStudy (anonymous):

\[5\sqrt{5x}\] I don't know..

OpenStudy (nory):

You add the 3 and the 15, then the 2x. The 2x can't be added to the 3 and the 15 because it has an x in it. So"-12√2x + 2√2x. Do you see how I got that?

OpenStudy (nory):

[Sorry--this is kind of a long, intensive math session...]

OpenStudy (anonymous):

I think you mixed it up. where did you get the \[2\sqrt{2x}\] ?

OpenStudy (nory):

Oops: 2x√2x. Thanks for catching my mistake.

OpenStudy (anonymous):

What would be the next step?

OpenStudy (nory):

There is no next step. There's nothing left to add, so you're done.

OpenStudy (anonymous):

\[-12\sqrt{2x}\] + \[2x \sqrt{2x}\] well for this problem, I have to find out what the person did wrong.

OpenStudy (nory):

I actually did the problem the way he did it the first time, and then I realized my mistake. What he did was he added the -12√2x and the 2x√2x, and you can't do that because the 2x√2x has an x in it (in front of the radical). It's like adding 2 and 5x^2: you can't do that, can you?

OpenStudy (anonymous):

Alright, thank you for all this information!(:

OpenStudy (nory):

Good work. And thanks for staying with me all this time. What was this, like an hour? :)

OpenStudy (anonymous):

Haha, 34 minutes :P